OpenStudy (anonymous):

Chef Pierre can do something unique. Using a secret process, he can bake a nearly perfectly spherical pie consisting of a vegetable filling inside a thick crust. The radius of the whole pie is 12 cm, and the radius of the filling is 8 cm. What is the volume of the crust alone, to the nearest unit? Use π ≈ 3.14.

OpenStudy (***[isuru]***):

ok! to do this ... u must volume of the pie - volume of the filling and a volume of a sphere is given by \[A = \frac{ 4 }{ 3 } \pi r^{3}\] so if r = radius of the pie and r note = radius of the filling then answer \[\frac{ 4 }{ 3 } \pi r^{3} - \frac{ 4 }{ 3 } \pi r _{o}^{3}\] now \[\frac{ 4 }{ 3 } \pi ( r^{3} - r_{o}^{3} )\] which is 4/3 x pi ( 12^3 - 8^3) 4/3 pi (1216 so the answer is 5090 . 98 cubic centimeters hope this will help ya! :)